# 6D Rigid Body Localization and Velocity Estimation via Gaussian Belief Propagation

### v2

@@ III-C Bivariate GaBP for Transformation Parameter Estimation
While the linear formulation and the message passing rules for the GaBP iterations are very similar to Algorithm 1, in the case of transformation parameter estimation based on equation (28), there exist two sets of variables $\theta_{k}$ with $k\in\{1,\ldots,K\}$ and $t_{\ell}$ with $\ell\in\{1,\ldots,K\}$, such that the GaBP rules are elaborated separately. To elaborate, first soft-IC is performed on the observed information respectively for the angle and translation variables as
@@
where the corresponding extrinsic means and variances are given by
@@
Finally the denoisers with a Gaussian prior are given by
@@
$\check{\psi}_{\theta:m,k}$ $=\frac{\phi_{\theta}\cdot\bar{v}_{\theta:m,k}^{[j]}}{\phi_{\theta}+\bar{v}_{\theta:m,k}^{[j]}}\in\mathbb{R},$ $\check{\psi}_{t:m,\ell}$ $=\frac{\phi_{t}\cdot\bar{v}_{t:m,\ell}^{[j]}}{\phi_{t}+\bar{v}_{t:m,\ell}^{[j]}}\in\mathbb{R},$ (44b)
@@
6:Initialise $\hat{\theta}_{m,k}^{[1]}$, $\hat{t}_{m,\ell}^{[1]}$, $\psi_{\theta:m,k}^{[1]}$, $\psi_{t:m,\ell}^{[1]}$;
@@
27:Obtain refined consensus estimates $\tilde{\theta}_{k}$ via eq. (LABEL:eq:del_t_final_est_theta);
Subsequently, the soft-replicas are iteratively updated similarly to equation (36), with a damping factor $\rho$, for $j_{\mathrm{max}}$ iterations of the message passing algorithm or a convergence criteria, after which the consensus estimates are obtained as
$\tilde{\theta}_{k}$ $=\bigg{(}\sum_{m=1}^{M}\frac{|h_{\theta:m,k}|^{2}}{\big{(}\sigma_{\theta:m,k}^{[j_{\mathrm{max}}]}\big{)}^{2}}\bigg{)}^{\!\!\!-1}\!\!\bigg{(}\sum_{m=1}^{M}\frac{h_{\theta:m,k}\cdot\tilde{z}_{\theta:m,k}^{[j_{\mathrm{max}}]}}{\big{(}\sigma_{\theta:m,k}^{[j_{\mathrm{max}}]}\big{)}^{2}}\bigg{)}\in\mathbb{R},$
$\tilde{t}_{\ell}$ $=\bigg{(}\sum_{m=1}^{M}\frac{|h_{t:m,\ell}|^{2}}{\big{(}\sigma_{t:m,\ell}^{[j_{\mathrm{max}}]}\big{)}^{2}}\bigg{)}^{\!\!\!-1}\!\bigg{(}\sum_{m=1}^{M}\frac{h_{t:m,\ell}\cdot\tilde{z}_{t:m,\ell}^{[j_{\mathrm{max}}]}}{\big{(}\sigma_{t:m,\ell}^{[j_{\mathrm{max}}]}\big{)}^{2}}\,\bigg{)}\in\mathbb{R}.$
While the message passing rules elaborated by equations (38)-(45) are complete to yield the estimated rotation angles and translaction vectors, due to the effective channel powers of $\bm{H}_{\theta}$ and $\bm{H}_{t}$ in equation (28) where the latter is typically much larger absolute positions of the anchors and sensors [^6] The effective channel powers are highly depedent on the sensor and anchor deployment structure, and for typical indoor sensing scenarios as illustrated in Figure 1, the anchor coordinates are of larger absolute value than the rigid body sensor coordinates, leading to the large power difference.. Such significant difference in effective channel powers lead to good estimation performance of the translation vector elements, but erroneous estimation performance of the rotation angles in a joint estimation described by the GaBP procedure.
In order to address the aforementioned error behavior of the rotation angle parameters $\bm{\theta}$, we propose an interference cancellation-based approach to remove the components corresponding to the translation of the sensors, and perform the GaBP again only on the rotation angle parameters. Namely, by using the estimated consensus translation vector $\tilde{\bm{t}}\triangleq[\tilde{t}_{1},\tilde{t}_{2},\tilde{t}_{3}]^{\mathsf{T}}\!\in\mathbb{R}^{3\times 1}$ obtained at the end of the GaBP via equation (LABEL:eq:del_t_final_est_t), the interference-cancelled system is given by
$\bm{z}_{n}^{\prime}\triangleq\bm{z}_{n}-\bm{H}_{t}\tilde{\bm{t}}=\bm{H}_{\theta}\bm{\theta}+\bm{\xi}_{n}\in\mathbb{R}^{M\times 1}.\vspace{-1ex}$ (46)
@@
$\tilde{z}{{}^{\prime}}_{\!\!\theta:m,k}^{[j]}=z^{\prime}_{m}-\sum_{i\neq k}h_{\theta:m,i}\hat{\theta}_{m,i}^{[j]}\in\mathbb{R},\vspace{-1ex}$ (47)
$\sigma_{\theta:m,k}^{2\,[j]}=\sum_{i\neq k}\big{|}h_{\theta:m,i}\big{|}^{2}\psi_{\theta:m,i}^{[j]}+N_{0}\in\mathbb{R},\vspace{-1ex}$ (48)
which is concatenated with the previously described bivariate GaBP to describe the complete estimation process of the rigid body transformation parameters $\bm{\theta}$ and $\bm{t}$, as summarized by Algorithm 2 above.
@@ IV-A Velocity Transformation Parameter-based System Model
Utilizing the same small-angle approximation [10] as before, the angular velocity matrix $[\bm{\omega}]^{\times}$ in equation (13) can be vectorized and decomposed into a linear system directly in terms of the individual velocities [11], yielding
@@ IV-B Linear GaBP for Sensor Velocity Estimation
It can be observed that the GaBP steps are equivalent to the ones presented in Section III-B. In what follows, we shall therefore be brief and described the steps adjusted to the new corresponding linear system of equations succintly.
@@
$\tilde{\dot{x}}_{k}=\bigg{(}\sum_{m=1}^{M}\frac{|\dot{g}_{m,k}|^{2}}{\sigma_{m,k}^{2[j_{\mathrm{max}}]}}\bigg{)}^{\!\!\!-1}\!\!\!\bigg{(}\sum_{m=1}^{M}\frac{\dot{g}_{m,k}\cdot\tilde{\dot{y}}_{m,k}^{[j_{\mathrm{max}}]}}{\sigma_{m,k}^{2[j_{\mathrm{max}}]}}\bigg{)}\in\mathbb{R}.\vspace{1ex}$ (60)
@@ IV-C Bivariate GaBP for Velocity Parameter Estimation
$\sigma_{\dot{t}:m,\ell}^{2\,[j]}$ $=\sum_{i=1}^{K}\big{|}b_{\omega:m,i}\big{|}^{2}\psi_{\omega:m,i}^{[j]}+\sum_{i\neq\ell}\big{|}b_{\dot{t}:m,i}\big{|}^{2}\psi_{\dot{t}:m,i}^{[j]},$ (63b)
@@
Finally, the denoisers with a Gaussian prior are given by
Using the estimated consensus translational velocity vector $\tilde{\bm{\dot{t}}}\triangleq[\tilde{\dot{t}}_{1},\tilde{\dot{t}}_{2},\tilde{\dot{t}}_{3}]^{\mathsf{T}}\!\in\mathbb{R}^{3\times 1}$ obtained at the end of the GaBP via equation (LABEL:eq:del_t_final_est_t), the interference-cancelled system is given by
$\bm{u}_{n}^{\prime}\triangleq\bm{u}_{n}-\bm{B}_{\dot{t}}\tilde{\bm{\dot{t}}}=\bm{B}_{\omega}\bm{\omega}+\bm{\zeta}_{n}\in\mathbb{R}^{M\times 1}.\vspace{-1ex}$ (70)
$\tilde{u}{{}^{\prime}}_{\!\!\omega:m,k}^{[j]}=u^{\prime}_{m}-\sum_{i\neq k}b_{\omega:m,i}\hat{\omega}_{m,i}^{[j]}\in\mathbb{R},\vspace{-1ex}$ (71)
@@
6:Initialise $\hat{\omega}_{m,k}^{[1]}$, $\hat{\dot{t}}_{m,\ell}^{[1]}$, $\psi_{\omega:m,k}^{[1]}$, $\psi_{\dot{t}:m,\ell}^{[1]}$;
@@ V-A RBL Parameter Estimation Performance Against the SotA
In this section, we present simulation results to demonstrate the effectiveness of the proposed multi-stage GaBP-based approach for RBL. Specifically, we compare the estimation performance of the two stages of the proposed approaches: a) the sensor position estimation via linear GaBP in Algorithm 1, and the RBL transformation parameter estimation via bivariate GaBP and interference cancellation-based refinement GaBP concatenated in Algorithm 2 for a stationary rigid body, and b) the sensor velocity estimation via Algorithm 3, and the RBL velocity parameter estimation via Algorithm 4 for a moving rigid body. The performance is compared against the relevant SotA RBL solution, where the time of arrival (TOA)-based sensor position estimation is performed via the approach in [29], the joint time difference of arrival (TDOA) and frequency difference of arrival (FDOA)-based sensor velocity estimation is performed via the approach in [30] and the RBL parameter estimation is performed via [11].
The simulation setup is the scenario illustrated in Figure 1, where the rigid body is composed of $N=8$ sensors positioned at the vertices of a unit cube at the origin, with sensor positions described by the conformation matrix given by
@@
The RBL rotation angles $\theta_{x},\theta_{y},\theta_{z}$ follow a zero-mean Gaussian distribution of variance $\phi_{\theta}=10$, and the RBL translation vector elements also follow a zero-mean Gaussian distribution of variance $\phi_{t}=5$. The RBL parameters in the moving scenario follow a zero-mean Gaussian distribution of variance $\phi_{\omega}=10$ for the angular velocity and a variance of $\phi_{\dot{t}}=5$ for the translational velocity.
@@
where $\hat{\bm{x}}^{[i]}$ is the RBL parameter vector (position, angle, or translation) estimated during the $i$-th Monte-Carlo simulation, $\bm{x}$ is the true RBL parameter vector, and $E=10^{4}$ is the total number of independent Monte-Carlo experiments used for the analysis, and is evaluated for different noise standard deviations $\sigma$ of equation (14) and (18), defined as $\sigma\!=\!\sigma_{w}\!=\!0.1\sigma_{\epsilon}$.
@@
Firstly, for the stationary scenario, Figure 3 illustrates the sensor position estimation performance of the proposed Algorithm 1 and the SotA closed-form solution of [29] for different noise standard deviations, also referred to as range error in meters. It can be observed that the proposed linear GaBP method outperforms the SotA approach for all range error regimes, which suggests that already the proposed preliminary positioning method can be used as an initializer for other SotA methods to perform RBL parameter extraction. In addition, Figures 5 and 5 illustrate the estimation performance of the rigid body translation vector and rotation angles respectively.
@@
Figure 4: RMSE of the proposed bivariate GaBP algorithm for rigid body translation estimation (Alg. 2) and the SotA method of [11], for various range errors $\sigma$.
@@
Figure 5: RMSE of the proposed bivariate GaBP algorithm for rigid body rotation angle estimation (Alg. 2) and the SotA method of [11], for various range errors $\sigma$.
@@
As mentioned in Section III, the translation parameter can be effectively estimated via the proposed bivariate GaBP of Algorithm 2 in light of the channel power difference effect, which is highlighted by the result of Figure 5. It can be seen that the proposed bivariate GaBP algorithm outperforms the SVD-based SotA approach [11] in terms of the rotation angle estimation, for all regimes of range error.
@@
Finally, Figure 5 illustrates the estimation performance of the rotation angles of the rigid body transformation. It can be seen that the estimation performance of the proposed concatenated GaBP also exhibits superiority to the SotA method, similarly to the behavior of the positioning and translation estimation performance. Due to the aforementioned channel power scaling effect which causes the noise power to become more prominent in the estimation of the variables for small range error regimes, the performance of the proposed method can be seen to approach the performance of the SotA in Figure 5. This trend is due to the common error floor usually seen in belief propagation algorithms at smaller error regions, which can be further optimized via procedures such as adaptive damping [41].
@@
Additionally, the ideal behavior of the proposed algorithm is illustrated via the matched filter bound (MFB) of the GaBP algorithm, which shows the expected superiority over all noise ranges, also illustrating that for high noise, the proposed method indeed reaches the optimal performance.
Next, the second set of results evaluates the performance of the proposed rigid body motion estimation algorithms within the moving rigid body scenario. First, Figure 7 illustrates the RMSE of the sensor velocity estimates of Algorithm 3 compared to the SotA solution of [30]. It can be observed that the proposed method outperforms the SotA in all velocity errors with a large performance gain.
@@
Figure 7: RMSE of the proposed bivariate GaBP algorithm for rigid body translational velocity estimation (Alg. 4) and the SotA method of [11], for various noise levels $\sigma$.
Figure 8: RMSE of the proposed bivariate GaBP algorithm for rigid body angular velocity estimation (Alg. 4) and the SotA method of [11], for various noise levels $\sigma$.
Finally, Figures 7 and 8 show the performance of the estimate velocity parameters, $i.e.$, the angular velocity and the translational velocity. In Figure 7, the translational velocity estimate of the proposed method in Algorithm 4 is compared to the SotA method of [11]. It can be observed that as previously, the proposed method outperforms the SotA for all range errors. In turn, Figure 8 illustrates the performance of the angular velocity estimation, where similar to the stationary rotation estimation results, the proposed method outperforms the SotA for all range errors, with a larger gain reaching the MFB for high noise and a smaller gain for low noise.
V-B Convergence Behavior and Complexity Analysis
Here, the performance of the proposed RBL parameter estimation algorithms are further assessed in terms of the computational complexity and convergence behavior, which is thoroughly investigated in [42, 43, 44], and compared against the SotA methods.
Figure 9: Convergence of the proposed algorithms for a noise level of $\sigma=10^{-1}$.

### v3

@@ III-C Bivariate GaBP for Transformation Parameter Estimation
While the linear formulation and the message passing rules for the GaBP iterations are very similar to Algorithm 1, in the case of transformation parameter estimation based on equation (28), there exist two sets of variables $\theta_{k}$ with $k\in\{1,\ldots,K\}$ and $t_{\ell}$ with $\ell\in\{1,\ldots,K\}$ as illustrated by the factor graph structure of Figure 4 such that the GaBP rules are elaborated separately. To elaborate, first soft-IC is performed on the observed information respectively for the angle and translation variables as
@@
where the extrinsic means and variances are given by
@@
Figure 4: Factor graph structure of the bivariate system in equation (28), where the variable nodes (circles) represent the unknown variables $\theta_{k}$ and $t_{\ell}$ and the factor nodes (squares) represent the transformed effective observations $z_{m}$. Edges are defined by the connecting entries of the effective channel matrices $\bm{H}_{\theta}$ and $\bm{H}_{t}$.
@@
Finally, the denoisers with a Gaussian prior are given by
@@
$\hskip-8.61108pt\check{\psi}_{\theta:m,k}=\frac{\phi_{\theta}\cdot\bar{v}_{\theta:m,k}^{[j]}}{\phi_{\theta}\!+\!\bar{v}_{\theta:m,k}^{[j]}}\in\mathbb{R},\;\check{\psi}_{t:m,\ell}=\frac{\phi_{t}\cdot\bar{v}_{t:m,\ell}^{[j]}}{\phi_{t}\!+\!\bar{v}_{t:m,\ell}^{[j]}}\in\mathbb{R},$ (45b)
@@
Subsequently, the soft-replicas are iteratively updated similarly to equation (37), with a damping factor $\rho$, for $j_{\mathrm{max}}$ iterations of the message passing algorithm or until a convergence criteria is met, after which the consensus estimates are obtained as
$\tilde{\theta}_{k}=\bigg(\sum\limits_{m=1}^{M}\frac{|h_{\theta:m,k}|^{2}}{\big(\sigma_{\theta:m,k}^{[j_{\mathrm{max}}]}\big)^{2}}\bigg)^{\!\!\!-1}\!\!\bigg(\sum\limits_{m=1}^{M}\frac{h_{\theta:m,k}\cdot\tilde{z}_{\theta:m,k}^{[j_{\mathrm{max}}]}}{\big(\sigma_{\theta:m,k}^{[j_{\mathrm{max}}]}\big)^{2}}\bigg)\in\mathbb{R},\;\;\;$ (46a)
$\tilde{t}_{\ell}=\bigg(\sum\limits_{m=1}^{M}\frac{|h_{t:m,\ell}|^{2}}{\big(\sigma_{t:m,\ell}^{[j_{\mathrm{max}}]}\big)^{2}}\bigg)^{\!\!\!-1}\!\bigg(\sum\limits_{m=1}^{M}\frac{h_{t:m,\ell}\cdot\tilde{z}_{t:m,\ell}^{[j_{\mathrm{max}}]}}{\big(\sigma_{t:m,\ell}^{[j_{\mathrm{max}}]}\big)^{2}}\,\bigg)\in\mathbb{R}.\;\;\;$ (46b)
While the message passing rules elaborated by equations (39)-(46) are complete to yield the estimated rotation angles and translation vectors, there is a difference in effective channel powers of $\bm{H}_{\theta}$ and $\bm{H}_{t}$ in equation (28), where the latter is typically much larger, due to the absolute positions of the anchors and sensors [^10] The effective channel powers are highly dependent on the sensor and anchor deployment structure, and for typical indoor sensing scenarios as illustrated in Figure 1, the anchor coordinates are of larger absolute value than the rigid body sensor coordinates, leading to the large power difference.. Such significant differences in effective channel powers lead to good estimation performance of the translation vector elements, but erroneous estimation performance of the rotation angles in a joint estimation described by the GaBP procedure.
@@
6:Initialize $\hat{\theta}_{m,k}^{[1]}$, $\hat{t}_{m,\ell}^{[1]}$, $\psi_{\theta:m,k}^{[1]}$, $\psi_{t:m,\ell}^{[1]}$;
@@
27:Obtain refined consensus estimates $\tilde{\theta}_{k}$ via eq. (46a);
In order to address the aforementioned error behavior of the rotation angle parameters $\bm{\theta}$, we propose an interference cancellation-based approach to remove the components corresponding to the translation of the sensors, and perform the GaBP again only on the rotation angle parameters. Namely, by using the estimated consensus translation vector $\tilde{\bm{t}}\triangleq[\tilde{t}_{1},\tilde{t}_{2},\tilde{t}_{3}]^{\mathsf{T}}\!\in\mathbb{R}^{3\times 1}$ obtained at the end of the GaBP via equation (46b), the interference-cancelled system is given by
$\bm{z}_{n}^{\prime}\triangleq\bm{z}_{n}-\bm{H}_{t}\tilde{\bm{t}}=\bm{H}_{\theta}\bm{\theta}+\bm{\xi}_{n}\in\mathbb{R}^{M\times 1}.$ (47)
@@
$\tilde{z}{{}^{\prime}}_{\!\!\theta:m,k}^{[j]}=z^{\prime}_{m}-\sum_{i\neq k}h_{\theta:m,i}\hat{\theta}_{m,i}^{[j]}\in\mathbb{R},$ (48)
$\sigma_{\theta:m,k}^{2\,[j]}=\sum_{i\neq k}\big|h_{\theta:m,i}\big|^{2}\psi_{\theta:m,i}^{[j]}+N_{0}\in\mathbb{R},$ (49)
which is concatenated with the previously described bivariate GaBP to describe the complete estimation process of the rigid body transformation parameters $\bm{\theta}$ and $\bm{t}$, as summarized by Algorithm 2.
@@ IV-A Velocity Transformation Parameter-based System Model
Utilizing the same small-angle approximation [11] as before, the angular velocity matrix $\bm{\omega}]^{\times}$ in equation (13) can be vectorized and decomposed into a linear system directly in terms of the individual velocities [12], yielding
@@ IV-B Linear GaBP for Sensor Velocity Estimation
It can be observed that the GaBP steps are equivalent to the ones presented in Section III-B. In what follows, we shall therefore be brief and describe the steps adjusted to the new corresponding linear system of equations succinctly.
@@
$\tilde{\dot{x}}_{k}=\bigg(\sum_{m=1}^{M}\frac{|\dot{g}_{m,k}|^{2}}{\sigma_{m,k}^{2[j_{\mathrm{max}}]}}\bigg)^{\!\!\!-1}\!\!\!\bigg(\sum_{m=1}^{M}\frac{\dot{g}_{m,k}\cdot\tilde{\dot{y}}_{m,k}^{[j_{\mathrm{max}}]}}{\sigma_{m,k}^{2[j_{\mathrm{max}}]}}\bigg)\in\mathbb{R}.\vskip 4.30554pt$ (61)
@@ IV-C Bivariate GaBP for Velocity Parameter Estimation
$\sigma_{\dot{t}:m,\ell}^{2\,[j]}$ $=\sum_{i=1}^{K}\big|b_{\omega:m,i}\big|^{2}\psi_{\omega:m,i}^{[j]}+\sum_{i\neq\ell}\big|b_{\dot{t}:m,i}\big|^{2}\psi_{\dot{t}:m,i}^{[j]}+N_{0},$
@@
6:Initialize $\hat{\omega}_{m,k}^{[1]}$, $\hat{\dot{t}}_{m,\ell}^{[1]}$, $\psi_{\omega:m,k}^{[1]}$, $\psi_{\dot{t}:m,\ell}^{[1]}$;
@@
Finally, the denoisers with a Gaussian prior are given by
Using the estimated consensus translational velocity vector $\tilde{\bm{\dot{t}}}\triangleq[\tilde{\dot{t}}_{1},\tilde{\dot{t}}_{2},\tilde{\dot{t}}_{3}]^{\mathsf{T}}\!\in\mathbb{R}^{3\times 1}$ obtained at the end of the GaBP via equation (46b), the interference-cancelled system is given by
$\bm{u}_{n}^{\prime}\triangleq\bm{u}_{n}-\bm{B}_{\dot{t}}\tilde{\bm{\dot{t}}}=\bm{B}_{\omega}\bm{\omega}+\bm{\zeta}_{n}\in\mathbb{R}^{M\times 1}.\vskip-4.30554pt$ (71)
$\tilde{u}{{}^{\prime}}_{\!\!\omega:m,k}^{[j]}=u^{\prime}_{m}-\sum_{i\neq k}b_{\omega:m,i}\hat{\omega}_{m,i}^{[j]}\in\mathbb{R},\vskip-4.30554pt$ (72)
@@ V-A RBL Parameter Estimation Performance Against the SotA
In this section, we present simulation results to demonstrate the effectiveness of the proposed multi-stage GaBP-based approach for RBL. Specifically, we compare the estimation performance of the two stages of the proposed approaches: a) the sensor position estimation via linear GaBP in Algorithm 1, and the RBL transformation parameter estimation via bivariate GaBP and interference cancellation-based refinement GaBP concatenated in Algorithm 2 for a stationary rigid body, and b) the sensor velocity estimation via Algorithm 3, and the RBL velocity parameter estimation via Algorithm 4 for a moving rigid body. The performance is evaluated in comparison with two relevant SotA RBL methods. The first method, presented in [12], involves a three-stage estimation procedure. Initially, sensor positions are estimated using a ToA-based approach as described in [31]. Subsequently, sensor velocities are jointly estimated from time difference of arrival (TDOA) and frequency difference of arrival (FDOA) measurements following the method in [32]. Finally, the RBL parameters are estimated using the approach outlined in [12].
The second SotA method considered is that proposed in [30], which is a high complexity and high accuracy technique. This method estimates the RBL parameters using semidefinite relaxation (SDR) applied to ToA and Doppler measurements.
The simulation setup is the scenario illustrated in Figure 1, where the rigid body is composed of $N=8$ sensors positioned at the vertices of a unit cube at the origin, with sensor positions described, in meters, by the conformation matrix given by
@@
The RBL rotation angles $\theta_{x},\theta_{y},\theta_{z}$ follow a zero-mean Gaussian distribution of variance $\phi_{\theta}=10{\color[rgb]{0,0,0}\text{deg}^{2}}$, and the RBL translation vector elements also follow a zero-mean Gaussian distribution of variance $\phi_{t}=5{\color[rgb]{0,0,0}\text{m}^{2}}$. The RBL parameters in the moving scenario follow a zero-mean Gaussian distribution of variance $\phi_{\omega}=10{\color[rgb]{0,0,0}(\text{deg}/\text{s})^{2}}$ for the angular velocity and a variance of $\phi_{\dot{t}}=5{\color[rgb]{0,0,0}(\text{m}/\text{s})^{2}}$ for the translational velocity.
@@
where $\hat{\bm{x}}^{[i]}$ is the RBL parameter vector (position, angle, or translation) estimated during the $i$-th Monte-Carlo simulation, $\bm{x}$ is the true RBL parameter vector, and $E=10^{4}$ is the total number of independent Monte-Carlo experiments used for the analysis, and is evaluated for different noise standard deviations [^11] In the moving scenario the range and Doppler noise is modeled coupled, such that a varying noise level effects both, the range and Doppler measurements. It is important to note that for certain scenarios a sensitivity analysis can become of interest, where the effect of the different types of noise are investigated individually, which is, however, not applicable to the proposed scenario. $\sigma$ of equation (14) and (18), defined as $\sigma\!\triangleq\!\sigma_{w}\!=\!0.1\sigma_{\epsilon}$. For the GaBP algorithms, the damping factor is set to $\rho=0.5$, and the maximum number of iterations is set to $j_{\mathrm{max}}=30$, which has empirically shown sufficient convergence.
@@
Firstly, for the stationary scenario, Figure 5 illustrates the sensor position estimation performance of the proposed Algorithm 1 and the SotA closed-form solution of [31] for different noise standard deviations, also referred to as range error in meters. It can be observed that the proposed linear GaBP method outperforms the SotA approach for all range error regimes, which suggests that already the proposed preliminary positioning method can be used as an initializer for other SotA methods to perform RBL parameter extraction. In addition, Figures 7 and 7 illustrate the estimation performance of the rigid body translation vector and rotation angles respectively.
@@
Figure 6: RMSE of the proposed bivariate GaBP algorithm for rigid body translation estimation (Alg. 2) and the SotA methods of [12] and [30], for various range errors $\sigma$.
@@
Figure 7: RMSE of the proposed bivariate GaBP algorithm for rigid body rotation angle estimation (Alg. 2) and the SotA methods of [12] and [30], for various range errors $\sigma$.
@@
As mentioned in Section III, the translation parameter can be effectively estimated via the proposed bivariate GaBP of Algorithm 2 in light of the channel power difference effect, which is highlighted by the result of Figure 7, even outperforming the high accuracy SDR-method of [30].
@@
Finally, Figure 7 illustrates the estimation performance of the rotation angles of the rigid body transformation. It can be seen that the estimation performance of the proposed concatenated GaBP also exhibits superiority to the SotA method of [12], similarly to the behavior of the positioning and translation estimation performance, but is slightly inferior to the SDR-method [30]. Due to the aforementioned channel power scaling effect which causes the noise power to become more prominent in the estimation of the variables for small range error regimes, the performance of the proposed method can be seen to approach the performance of the 2 Stage SotA in Figure 7. This is due to error floors usually seen in belief propagation algorithms at smaller error regions, which can be further optimized via adaptive damping [44].
@@
Additionally, the ideal behavior of the proposed algorithm is illustrated via the matched filter bound (MFB) of the GaBP algorithm, where the algorithm soft-replicas are initialized with the true values [45, 46] to illustrate the optimal performance bound, which shows the expected superiority over all noise ranges, also illustrating that for high noise, the proposed method indeed reaches the optimal performance.
@@
Figure 9: RMSE of the proposed bivariate GaBP algorithm for rigid body translational velocity estimation (Alg. 4) and the SotA methods of [12] and [30], for various noise levels $\sigma$.
Figure 10: RMSE of the proposed bivariate GaBP algorithm for rigid body angular velocity estimation (Alg. 4) and the SotA methods of [12] and [30], for various noise levels $\sigma$.
Furthermore, it should be noted that the use of the small angle approximation, although valid for the scenario considered in this work, may lead to singularities when the rotation angles become large [^12] This problem can be mitigated by replacing the linear approximation with a quadratic approach, combined with a strategy whereby the variable squared is treated as another variable correlated to that latter [47]. To this end, however, a non-trivial redesign of the GaBP algorithm is required, which is out of the scope of this article and will be pursued in a follow up work.. A notable example is the gimbal lock phenomenon, in which two rotation axes align, causing a loss of one degree of freedom in the parameterization.
Even before such issues occur, the Jacobian matrix may become ill conditioned, resulting in numerical instability during the estimation process. While such issues do not manifest in the current simulation setup, scenarios involving high dynamics or large rotation angles may necessitate the adoption of alternative estimation techniques. Approaches based on quaternions [48] or Lie groups [49] have been proposed to address these challenges, offering robust estimation without susceptibility to singularities.
Finally, it is observed that, even in cases where the rotation angles are large and the small angle approximation is no longer valid, the estimation performance remains comparable to that of smaller angle scenarios under conditions of high range error. However, in low range error regimes, the performance converges to an error floor due to the limitations imposed by the small angle approximation.